# Uncertainty Comes for Free: Human-in-the-Loop Policies with Diffusion Models

### v1

@@ Abstract
Human-in-the-loop (HitL) robot deployment has gained significant attention in both academia and industry as a semi-autonomous paradigm that enables human operators to intervene and adjust robot behaviors at deployment time, improving success rates. However, continuous human monitoring and intervention can be highly labor-intensive and impractical when deploying a large number of robots. To address this limitation, we propose a method that allows diffusion policies to actively seek human assistance only when necessary, reducing reliance on constant human oversight. To achieve this, we leverage the generative process of diffusion policies to compute an uncertainty-based metric based on which the autonomous agent can decide to request operator assistance at deployment time, without requiring any operator interaction during training. Additionally, we show that the same method can be used for efficient data collection for fine-tuning diffusion policies in order to improve their autonomous performance. Experimental results from simulated and real-world environments demonstrate that our approach enhances policy performance during deployment for a variety of scenarios.
Index Terms:
Human-in-the-loop policies, policy fine-tuning.
11footnotetext: Equal contributions.
I Introduction
@@
Human-in-the-Loop (HitL) robot deployment is a paradigm where a human operator can intervene and assist the robot during deployment. This paradigm is seeing increasing adoption in cases where robots must continue to operate adequately even in corner cases not foreseen before deployment.
In parallel, even as recent advances in policy learning have shown significant improvements in robustness at deployment time, current methods can still fail due to problems such as data distribution shift or incomplete state observability. This issue is particularly pronounced in robotics, where available datasets are far smaller compared to other domains.
Figure 1: Human-in-the-loop policies: We assume a human-in-the-loop robotic agent that actively requests human assistance when necessary. In our method, the robot uses an underlying diffusion policy for autonomous operation. During deployment, we leverage the denoising process inherent to diffusion models in order to identify states with high uncertainty, where the robot seeks operator intervention. We show that this approach allows for task improvement with a small number of such operator calls. Furthermore, the human-operated data can also be used to fine-tune the policy, improving the robot’s autonomous execution performance in future deployment.
@@
To address this issue, HitL methods can be a natural fit for learning agents. The robot can operate autonomously when possible, leveraging the ability of policy learning to execute complex motor control tasks. An expert operator can take over to deal with corner cases, ensuring overall task success. However, deploying human-in-the-loop policies can be labor-intensive if it implies constant monitoring of the robot’s behavior by the human operator, or frequent, interruptive interventions. Such limitations can make HitL deployment both impractical and costly in many applications.
In this work, we propose a data-driven approach for generating HitL policies. Our method is based on diffusion models, and does not require any additional computation or human intervention during training. Instead, we utilize the intrinsic properties of diffusion models, specifically the denoising process, to allow the agent to compute its internal uncertainty during deployment. When this uncertainty exceeds a certain threshold, the agent proactively seeks human assistance. Furthermore, a key feature of our method is its dual utility: it can be used not only for HitL deployment, but also for collecting additional data to fine-tune the policy.
We test our method on experimental setups where uncertainty can arise from partial observability (where our uncertainty metrics can enable effective supervision), data distribution shifts (where targeted data collection can also improve policy performance) and action multi-modality (where human guidance can help steer the robot into states where mode selection becomes unambiguous). Our main contributions are as follows:
@@
  • To the best of our knowledge, we are the first to propose using uncertainty estimation with diffusion policies for efficient HitL deployment of a diffusion-based learning agent. Furthermore, our approach does not require human-robot interactions during training, and incurs minimal computational overhead during deployment.
@@
  • We validate our method across multiple types of deployment challenges, in both simulated and real environment. Experimental results show that our approach requires fewer human interventions to and achieves higher task performance compares to alternative learning-based HitL agents.
@@
We see our work as complementary to approaches aiming to improve the robustness of fully autonomous learning-based agents. Until corner cases can be fully eliminated and such agents can generalize to any situation that can be encountered during deployment, HitL methods have the potential to bridge the gap to applications, and help collect critical training data for further policy improvements. We believe this work is a step in this direction.
II Related Work
Despite recent advances in generalist robot manipulation policies [8, 25], particularly in domestic settings [20, 36], deploying these policies in real-world scenarios remains a significant challenge. Robots often face substantial data distribution shifts when introduced to new environments. In SIMPLER [17], Li et al. show the sensitivity of policies to minor changes, such as variations in robot arm texture, which can cause performance drops exceeding 20%. Additionally, real-world deployment is hampered by incomplete and noisy observations due to environmental factors such as lighting conditions, occlusions, clutter, and weak texture features [9]. The inherent complexity of modeling multi-modal distributions in human demonstrations further exacerbates these issues, particularly in behavior cloning, where stochastic sampling and initialization procedures introduce additional challenges [7, 10, 23, 29]. Our work addresses these deployment difficulties by incorporating human assistance at critical moments. By identifying high-uncertainty states during deployment, our method facilitates timely human intervention, enabling robots to overcome distribution shifts, handle incomplete observations, and navigate complex multi-modal action spaces more effectively.
Human-in-the-loop approaches have been widely explored to improve robot manipulation policies by leveraging human feedback in various forms, including interventions [23, 31], preferences [16], rankings [4], scalar feedback [22], and human gaze [38]. Recent works [19, 21] have demonstrated the effectiveness of human assistance. For example, HIL-SERL achieves state-of-the-art performance in vision-based real-world reinforcement learning, while Sirius integrates human trust signals to enhance behavioral cloning. The HitL systems is also extensively applied in autonomous driving. For example re-routing systems from ZOOX [24] enable human operators to intervene when vehicles encounter challenging scenarios.
However, these approaches often focus on incorporating human feedback during training while neglecting the critical question of when human assistance is most necessary during deployment. They typically require continuous human supervision, which is inefficient and impractical in real-world settings. In contrast, our work introduces an uncertainty-aware diffusion model that actively identifies deployment scenarios where human expertise is most beneficial. By requesting assistance only during high-uncertainty states, we enable a more efficient and scalable HitL framework that minimizes the reliance on constant human oversight. The closest work to ours is HULA [30], a reinforcement learning approach that explicitly models the uncertainty of an agent. However, HULA has limited application in the setting of online reinforcement learning, while our works focus on deriving HitL policies from offline dataset.
In this work, we leverage diffusion models to produce HitL policies. Diffusion models have recently achieved remarkable success in robotics and decision-making tasks [2, 7, 26, 28, 32, 34, 37]. These models are particularly adept at capturing complex behaviors and multi-modal trajectory distributions when trained on high-quality demonstration data. However, collecting perfect demonstration datasets is often infeasible due to limitations in data quality and the prevalence of suboptimal demonstrations. To address this, researchers have proposed methods such as guiding the denoising process with external objectives like reward signals or goal conditioning [1, 5, 13, 18, 33]. Other strategies include integrating Q-learning and weighted regression in offline [6, 35] or online [11, 14, 27] settings. Our work takes a novel approach by leveraging the inherent distribution modeling capabilities of diffusion models to quantify uncertainty in action modes for each state. This allows us to identify high-uncertainty states where human intervention is most impactful. By targeting these critical states, we create a more efficient and focused human-in-the-loop policy deployment framework, enhancing both performance and scalability.
III Method
We start from the basic premise of HitL robotic agents: the robot is expected to generally act autonomously, but a human operator is available to provide teleoperation commands should the robot require them. Our method is designed to determine when the agent should request such expert assistance, making effective use of a limited number of such calls during deployment. Furthermore, since we use a learning agent, we aim to eliminate the need for expert intervention during the training phase, as that would likely place a large burden on the operator. This means that the agent has no quantitative knowledge about the effect of assistance, except for the assumption that it effective for task completion.
To achieve this, we utilize diffusion models [7] as our policy class. Diffusion policies offer two key advantages: (1) they have demonstrated robust performance in imitation learning tasks, and (2) their generative process involves an iterative denoising mechanism which we can leverage for insight into the agent’s decision-making process. Specifically, we use the denoising information to compute an uncertainty metric for the policy, which is then used during deployment to determine when human intervention is most beneficial. Importantly, the policy is trained using an offline dataset and does not rely on human assistance during the training phase. Finally, we show that data collected during the operator interventions can be incorporated back into training through a fine tuning process that further improves policy performance.
In this section, we first re-iterate general diffusion policies concepts, describe how our method utilizes their generative process to compute an uncertainty metric, and finally discuss how this metric is applied to improve policy deployment performance.
III-A Background: diffusion policies
@@
Diffusion policies generate actions through an action-denoising process, leveraging denoising diffusion probabilistic models (DDPM). A DDPM models a continuous data distribution $p(x^{0})$ as reversing a forward noising process from $x^{0}$ to $x^{K}$, which is defined as a Markovian chain with Gaussian transition. Here, $x^{K}$ represents a sample of $\mathcal{N}(0,\sigma^{2}\mathcal{I})$. To generate actions, the model learns to predict the transition (i.e. noises added during each step) and map the data from $x^{K}$ back to $x^{0}$. Sampling begins with a random Gaussian noise and iteratively refines it to produce a denoised output.
Specifically, the generative process of a diffusion policy $\pi(\mathrm{A}|\mathrm{O})$, where $\mathrm{O}$ and $\mathrm{A}$ are observations and actions, starts by sampling a random noise $a^{K}_{t}$ and iteratively remove noises by:
$a^{k-1}_{t}=\beta(a^{k}_{t}-\gamma\epsilon_{\theta}(o_{t},a^{k}_{t},k)+\mathcal{N}(0,\sigma^{2}\mathcal{I})),$
where $\beta$, $\sigma$ and $\gamma$ are functions of iteration step $k$. To train a diffusion policy, we learn a score function $\epsilon_{\theta}$ that predicts the noises used during forward noising:
$\centering\mathcal{L}=MSE(\epsilon^{k},\epsilon_{\theta}(o_{t},a^{k}_{t}+\epsilon^{k},k))\@add@centering$ (1)
III-B Denoising-based uncertainty metric
@@
To estimate the uncertainty of a diffusion-based agent, our method leverages the generative process inherent in the diffusion policy. In particular, we note that, when operating in task-space control (where the action space consists of desired end-effector poses), the noise predicted during the generative process can be interpreted as a vector field pointing toward the target distribution in the next state. We can thus leverage this vector field to analyze whether a diffusion-based agent is confident about its generative target.
We assume that our policy is operating on task space control and a diffusion policy outputs absolute end-effector (e.e.) poses and manipulator state. We also assume that the current e.e. pose is available as an input for action denoising. Our goal is thus to estimate an uncertainty metric $\texttt{Uncertainty}(o_{t})$ where $o_{t}$ is the observation at time step $t$.
To compute this metric, we first sample a set of points $A_{sampled}$ within a distance $r$ from the current pose. For each sampled point, we feed the diffusion policy forward and predict the noise vectors required to sample actions:
$v=\epsilon_{\theta}(o_{t},a_{sampled},0)$ (2)
These predicted noise vectors represent the directions toward the data distribution that the policy aims to recover. In this work, we use these denoising vectors to estimate uncertainty, defined as $\texttt{Uncertainty}(o_{t})=f(V_{t})$, where $V_{t}$ represents the set of denoising vectors at time step $t$.
@@
A critical characteristic of diffusion policies is their ability to capture multi-modality in the underlying human demonstrations (starting from a given state, there might be multiple distinct action trajectories that accomplish the desired task). Thus, at any given moment, the denoising vector field could reflect the multi-modal nature of the data, and naive variance estimation of the vector field may fail to capture this effect.
$\max_{n,\theta_{\mathrm{gmm}}}P(V_{t};n,\theta_{\mathrm{gmm}}),$
where $n$ is the number of modes and $\theta_{\mathrm{gmm}}$ is the parameters of a Gaussian mixture model. With the best-fit GMM, we then can estimate an agent’s uncertainty at state $s$. We first evaluate the divergence between each mode:
$D(V_{t})=\frac{1}{n(n-1)}\sum_{i,j}1-S_{c}(g_{i},g_{j})$
where,
$S_{c}(g_{i},g_{j})=\frac{g_{i}\cdot g_{j}}{||g_{i}||\cdot||g_{j}||}$
and $g_{i}$ represents the mean of the $i^{th}$ mode and $S_{c}$ represents cosine similarity between two vectors. We also evaluate the GMM variance as part of the uncertainty estimation:
$\texttt{Var}_{g}(V_{t})=\sum_{i}p(v_{i})\texttt{Var}(v_{i})$ (3)
where Var represents the variance of vector data. Putting them together, we can estimate the overall uncertainty as:
$\texttt{Uncertainty}(o_{t})=D(V_{t})+\alpha\texttt{Var}_{g}(V_{t}),$ (4)
III-C Uncertainty-based intervention and policy fine-tuning
@@
2:     Sample a set of points uniformly within the radius of $r$
3:     Feed forward the diffusion policy $\pi$ to collect a set of denoising vectors $V_{t}$
4:     Estimate uncertainty in this state using Eq.4
@@
8:     else
9:        Execute an action $a_{t}$ from the policy $\pi(a_{t}|o_{t})$
10:     end if
11:  end while
12:  if fine-tune then
13:     while fine-tuning not done do
14:        Sample a batch of data from $\mathcal{D}_{ft}$ and $\mathcal{D}_{train}$
15:        Optimize $\pi$ using Eq. 1
16:     end while
17:  end if
@@
Putting all components together, the final pipeline contains three main steps: 1. train a diffusion policy; 2. deploy the policy, and request operator control whenever policy uncertainty estimated by our metric exceeds a pre-set threshold; 3. (optional) use the human intervention data to fine-tune the diffusion policy.
To envision the intended applicability of this framework, let us consider three types of deployment issues that typically cause uncertainty for learning-based agents:
  • Case 1: Data distribution shift. For example, visual observation distribution shift can be caused by change of lighting condition. A special case for robotics is the change of dynamics that is caused by interacting with novel objects.
  • Case 2: Incomplete state observability. The common approach to address this can include redesigning, adding or moving sensors. However, in the general case, it is very difficult to have a sensor setup that guarantees full observations for all the tasks.
  • Case 3: Incorrect choice between different action modes. In cases where the agent is presented with a discrete choice between two possible action trajectory modes equally represented during training, diffusion policies are naturally well-equipped to make such choices. However, it is possible that, due to task under-specification, the choice will be incorrect for the given goal.
During policy execution, these problems may not be present in all states. In fact, many states are easy to make decisions. For example, moving the arm in free space is often easy and does not require human attention to help the robot. Our first goal is to identify when the issues arise, and request help at that time. We posit that, by combining uncertainty due to variance within a mode with uncertainty due to presence of multiple modes, our framework is positioned to recognize all three cases above. Furthermore, we believe that a HitL framework is particularly well equipped to handle Case 3, since a few steps under teleoperator control can “steer” the policy towards the desired mode. Case 1 lends itself well to fine-tuning based on the novel data collected during teleoperation. Finally, Case 2 is the most difficult, since, by definition, correct decision making is impossible without changing the available observation. We design our experiment set to test a range of scenarios covering these situations.
IV Experiments

### v2

@@ Abstract
Human-in-the-loop robot deployment has gained significant attention in both academia and industry as a semi-autonomous paradigm that enables human operators to intervene and adjust robot behaviors at deployment time, improving success rates. However, continuous human monitoring and intervention can be highly labor-intensive and impractical when deploying a large number of robots. To address this limitation, we propose a method that allows diffusion policies to actively seek human assistance only when necessary, reducing reliance on constant human oversight. To achieve this, we leverage the generative process of diffusion policies to compute an uncertainty-based metric based on which the autonomous agent can decide to request operator assistance at deployment time, without requiring any operator interaction during training. Additionally, we show that the same method can be used for efficient data collection for fine-tuning diffusion policies in order to improve their autonomous performance. Experimental results from simulated and real-world environments demonstrate that our approach enhances policy performance during deployment for a variety of scenarios.
Keywords: Human-in-the-loop Policies, Policy fine-tuning
1 Introduction
@@
Human-in-the-Loop (HitL) operation is a paradigm where a human operator can intervene and assist a robot during deployment. This paradigm is seeing increasing adoption in cases where robots must continue to operate adequately even in corner cases not foreseen before deployment.
@@
In parallel, even as recent advances in policy learning have shown significant improvements in robustness at deployment time [1, 2, 3, 4], current methods can still fail due to problems such as data distribution shift [5] or incomplete state observability [6]. To address this issue, HitL methods can be a natural fit for learning agents: the robot operates autonomously when possible, leveraging the ability of policy learning to execute complex motor control tasks. An expert operator can take over for corner cases, ensuring task success. However, deploying HitL can be labor-intensive and impractical if it implies constant monitoring of the robot’s behavior by the human operator, or frequent interventions.
In this work, we propose a data-driven approach for generating HitL policies. We start from the basic HitL premise: the robot generally acts autonomously, but a human operator is available to provide teleoperation commands should the robot require them. Our method is designed to determine when the agent should request such expert assistance, making effective use of a limited number of such calls during deployment. We also remove the need for expert intervention during the training phase, as that would place a large burden on the operator. This means that the agent has no knowledge about the effect of assistance, except for the assumption that is effective for task completion.
To achieve this, we utilize diffusion models [1] as our policy class. Diffusion policies offer two key advantages: (1) they have demonstrated robust performance in imitation learning tasks, and (2) their generative process involves an iterative denoising mechanism which we can leverage for insight into the agent’s decision-making process. Specifically, we use the denoising information to compute an uncertainty metric for the policy, which is then used during deployment to determine when human intervention is most beneficial (Fig. 1). To achieve this, we directly leverage the noise prediction model learned during the policy training process. Therefore, uncertainty estimation does not require training any additional models, imposes a minimal cost at run time, and can thus be considered a “free” byproduct of diffusion policy training. Finally, we show that data collected during the operator interventions can be incorporated back into training through a fine-tuning process that further improves policy performance. Our main contributions are as follows:
  • $\bullet$
@@
    We propose a novel method for HitL policy execution using uncertainty estimation with diffusion policies. Our method does not require human-robot interaction during training, and incurs minimal computational overhead during deployment.
  • $\bullet$
@@
    We validate our method across multiple types of deployment challenges, in both simulated and real environments. Experimental results show that our approach requires fewer human interventions and achieves higher task performance compared to alternative learning-based HitL agents.
  • $\bullet$
@@
Figure 1: HitL policies with denoising uncertainty: We propose using denoising uncertainty as a metric for deciding when to request (human) expert assistance. Predicted de-noising vectors in end-effector position space (illustrated here via arrows on end-effector position) are collected in a vector field, whose inter-mode divergence and intra-mode variance are used to compute policy uncertainty; when this measure exceeds a threshold, operator assistance is requested. We also show that ensuing teleoperation data can be used to fine-tune policies, achieving notable performance improvements with minimal additional data.
2 Related Work
Human-in-the-loop (HitL) approaches have been widely explored to enhance robot manipulation by incorporating various human feedback modalities, such as interventions [7, 8], preferences [9], rankings [10], scalar feedback [11], and gaze [12]. Recent works like HIL-SERL [13] and Sirius [14] show strong performance by leveraging human input during training. HitL methods are also common in autonomous driving, e.g., ZOOX’s re-routing system [15] allows operators to assist in challenging scenarios. However, most approaches rely on continuous human supervision and monitoring when assistance is actually needed during deployment. We propose an uncertainty-aware diffusion model that selectively requests human help in high-uncertainty states, reducing the need for constant oversight. Unlike HULA [16], which models uncertainty for online RL, our approach derives HitL policies from offline data, enabling more efficient and scalable deployment.
Our work is closely related to interactive imitation learning (IIL), where a learning agent queries an expert for additional labels during policy execution and augments the training dataset with expert demonstrations [17]. An effective data collection strategy for IIL is human-gated DAgger [18], which relies on a human to continuously monitor and intervene during robot execution. However, such continuous supervision is inefficient and undermines the goal of robot learning. To make HitL systems practical, the robot must be strategic in when and how it requests human assistance. Prior work has explored budgeting expert queries in IIL. For example, Hoque et al. reduce human effort by constraining robot queries using models of limited human attention [19] or fixed intervention budgets [20]. However, these methods require extra training procedures during training or deployment, increasing computational demands. This overhead is especially problematic during policy execution, where it may introduce additional latency. In contrast, our approach avoids additional training and supports efficient parallelization during deployment, resulting in minimal runtime overhead. Other approaches leverage action consistency [21], diffusion loss [22], or online conformal prediction [23] to decide when to query the human. One crucial aspect of this line of work is the selection or online-tuning of their thresholds. To minimize human efforts in a HitL pipeline, our method leverages the multi-modal nature of human demonstrations to detect critical states where the robot is uncertain. We demonstrate that our method is robust to threshold selection in different scenarios.
3 Method
We begin the description of our method with a short recap of diffusion policies, specifically the action denoising process which we will then leverage to introduce our metric for quantifying uncertainty. After that, we describe how this metric can be used by a HitL diffusion policy agent to determine when to request operator assistance, and finally how teleoperation data obtained through this method can in turn be used to fine-tune the original policy.
@@ 3 Method
Diffusion policies generate actions through an action-denoising process, leveraging denoising diffusion probabilistic models (DDPM). A DDPM models a continuous data distribution $p(a^{0})$ as reversing a forward noising process from $a^{0}$ to $a^{K}$, where $a^{K}$ is Gaussian noise sampled from $\mathcal{N}(0,\sigma^{2}\mathcal{I})$. The generative process $\pi(\mathrm{a_{t}}|\mathrm{o_{t}})$, where $\mathrm{o_{t}}$ and $\mathrm{a_{t}}$ are robot observations and actions at time step $t$, starts by sampling an action $a^{K}_{t}$ as random noise, and then iteratively denoises it using:
$a^{k-1}_{t}=\beta(a^{k}_{t}-\gamma\epsilon_{\theta}(o_{t},a^{k}_{t},k)+\mathcal{N}(0,\sigma^{2}\mathcal{I}))$ (1)
where $\beta$, $\sigma$ and $\gamma$ are functions of iteration step $k$. $\epsilon_{\theta}(o,a^{k},k)$ is a learned model that predicts the noise to be removed at each denoising step.
3.1 Denoising-based uncertainty metric
@@
To estimate the uncertainty of a diffusion-based agent, our method leverages the generative process described above. In particular, we assume that our diffusion policy is operating on task space control, which is a very common case in recent diffusion-based robot policy learning methods [24, 25, 26], and outputs absolute end-effector poses as part of its action vector. In this case, the noise predicted (and removed) during the generative process can be interpreted as a vector field pointing toward the distribution for intended end-effector pose the at the next step. We can thus leverage this vector field to analyze whether the diffusion-based agent is confident about its generative target.
Our goal is to estimate an uncertainty metric $\texttt{Uncertainty}(o_{t})$ where $o_{t}$ is the observation at time step $t$. We begin by sampling a set of end-effector poses $\mathcal{A}^{s}_{t}$, where each entry $a^{s}_{t}\in\mathcal{A}^{s}_{t}$ is within a distance $r$ from the current pose. When operating in task-space position control, each of these samples can be interpreted as an action vector. We can thus feed these samples through the diffusion policy noise prediction model, and collect the predicted noise vectors: let the set $\mathcal{V}^{s}_{t}$ contain all vectors $v^{s}_{t}=\epsilon_{\theta}(o_{t},a^{s}_{t},0)$ computed for each $a^{s}_{t}\in\mathcal{A}^{s}_{t}$. This vector field encodes directions toward the action distribution that the policy aims to recover. We will use these denoising vectors to estimate uncertainty, defined as $\texttt{Uncertainty}(o_{t})=f(\mathcal{V}^{s}_{t})$.
@@
The simplest method to assess uncertainty is to consider the variance of the vector field $\mathcal{V}^{s}_{t}$. However, diffusion policies are often used for their ability to capture multi-modality in the underlying demonstrations: from any given state, there might be multiple distinct action trajectories that accomplish the desired task. Thus, the denoising vector field could reflect the multi-modal nature of the demonstration data, and naive variance estimation of the vector field may fail to capture this effect.
@@
2:     Sample set $\mathcal{A}^{s}_{t}$ of poses within radius $r$ of end-effector
3:     Compute vector field $\mathcal{V}^{s}_{t}$ containing $v^{s}_{t}=\epsilon_{\theta}(o_{t},a^{s}_{t},0)$ for each $a^{s}_{t}\in\mathcal{A}^{s}_{t}$
4:     Estimate policy uncertainty based on $\mathcal{V}^{s}_{t}$ using Eq. (3)
@@
8:         else: Execute action $a_{t}$ from the policy $\pi(a_{t}|o_{t})$
9:     end if
10:  end while
11:  if fine-tune then
12:     while fine-tuning not done do
13:        Sample a batch of data from $\mathcal{D}_{ft}$ and $\mathcal{D}_{train}$
14:        Optimize $\pi$ with sampled data
15:     end while
16:  end if
$\max_{n,\theta_{\mathrm{g}}}P(\mathcal{V}^{s}_{t};n,\theta_{\mathrm{g}}),$
where $n$ is the number of modes and $\theta_{\mathrm{g}}$ contains the parameters of the GMM. With the best-fit GMM, we then estimate the agent’s uncertainty. We first evaluate the divergence between each mode:
$D(\mathcal{V}^{s}_{t})=\frac{1}{n(n-1)}\sum_{i,j}1-S_{c}(g_{i},g_{j})\text{, where }S_{c}(g_{i},g_{j})=\frac{g_{i}\cdot g_{j}}{||g_{i}||\cdot||g_{j}||}$
Here, $g_{i}$ represents the mean of the $i^{th}$ mode and $S_{c}$ represents cosine similarity between two vectors. We also evaluate the GMM variance as part of the uncertainty estimation:
$\texttt{Var}_{g}(\mathcal{V}^{s}_{t})=\sum_{i}p(v_{i})\texttt{Var}(v_{i})$ (2)
where Var represents the variance of vector data and $v_{i}$ represents vector samples belongs to the $i^{th}$ mode of the GMM. Putting them together, we can estimate the overall uncertainty as:
$\texttt{Uncertainty}(o_{t})=D(\mathcal{V}^{s}_{t})+\alpha\texttt{Var}_{g}(\mathcal{V}^{s}_{t}),$ (3)
3.2 Uncertainty-based intervention and policy fine-tuning
@@
Putting all components together, the method contains three main steps: 1. train a diffusion policy; 2. deploy the policy, and request operator control if policy uncertainty estimated by our metric exceeds a preset threshold; 3. (optional) use human intervention data to fine-tune the diffusion policy.
4 Experiments